OpenStudy (bananas1234):

2. A prism has total surface area of 158 m2 and volume of 120 m3. If the length, width, and height are reduced to one third their original sizes, what will be the following? Round to tenths if necessary (a) the new surface area (b) the new volume

OpenStudy (anonymous):

(a) the new surface area

OpenStudy (bananas1234):

Thanks, but i have to show my work, can you help with that? @EpicGammer

OpenStudy (welshfella):

(b) take a simple example suppose it was a simple cube with each side = 1 cm and volume = 1*1*1 = 1 cm3 each side will be reduced to 1/3 cm so new volume = 1/3 8 1/3 * 1/3 = 1/27

OpenStudy (welshfella):

* that 8 is supposed to be a *

OpenStudy (welshfella):

It will be the same ratio for your prism of volume 120 m3

OpenStudy (welshfella):

120 * 1/27